# Design and Implementation of a Scalable Financial Exchange in the Public Cloud

### v8

@@ 4.4 Remarks on Multicast Packet Losses
In on-premises exchanges, the multicast service does not provide reliable delivery. Instead, a separate "rewinder" service is deployed that handles any retransmission requests from MPs [25, 26]. MPs use sequence numbers of the messages to detect gaps and request a retransmission. The exchange infrastructure is engineered to provide negligible packet losses so the MPs only have to contact the rewinder rarely.
The Jasper multicast service needs to similarly provide low losses. Our hedging techniques help lower the packet losses, but we observe that even without our hedging, the losses in outbound direction are small: we observe 0.007% lost packets in a benchmark on GCP while exhausting 80% of egress bandwidth (more details are in Appendix H).
5 Inbound Communication
The matching engine (ME) is located at the root VM of the tree, while orders come from leaf nodes (MPs/traders). Using the proxy tree in reverse for order submission helps reduce the submission load on the ME. Intermediate proxies gather orders from multiple downstream nodes and forward them upstream, so the root VM simultaneously processes packets from only proxies instead of MPs. The reduced fan-in decreases packet losses. In a benchmark with 10 VMs each submitting 5K orders per second to the ME, the packet loss rate is 1.58%. Adding two intermediate proxies to create a tree structure lowers this loss rate to 0.68%, as the proxies alleviate incast congestion by relaying orders to the ME.
A tree alleviates some incast congestion, but is not sufficient especially when order bursts occur. We use packet pacing at the proxies to tackle such bursts. Avoiding drops by pacing the packets comes at the cost of filling up queues at the proxies which may increase order submission latency. To maintain low latency for orders and achieve high order matching rate of ME, we design a distributed packet scheduling policy, implemented at the proxies. The policy, realized through a priority queue, Limit Order Queue (LOQ), prioritizes orders that have prices closest to the current mid-price.
5.1 Pacing Outgoing Packets
Proxies continuously request the OWD estimates of packets sent to parent proxies or the ME. An increasing OWD is an indicator of congestion at the receiver (i.e., parent proxy or ME) and the proxies use this to slow down – increase the interpacket duration – to avoid packet drops.
Aggressive pacing can lead to bandwidth underutilization, while insufficient pacing risks excessive packet drops due to congestion. Ideal pacing balances these concerns by optimizing bandwidth usage, and adjusting pacing to avoid drops during bursts while reclaiming bandwidth once bursts vanish. In Jasper, prioritizing drop prevention is more important than operating at the maximum bandwidth. Sharding techniques for MEs can improve throughput [4], hence our focus is on minimizing packet drops during bursts. Nonetheless, it is favorable to utilize as much of the throughput as possible if it does not introduce substantial packet loss.
Our pacing heuristic calculates an amount of time to wait before sending the next packet (the pacing quantum) based on how much the most recent OWD estimate () deviates from a threshold OWD . is computed as the median OWD of all the packets sent in the last . The heuristic increases the pacing quantum by if increases by over . And the quantum is decreased by if falls below by . The heuristic is aggressive in increasing the pacing as compared to when decreasing the pacing because it is critical to avoid packet drops. When is sensed to be decreasing below , we slowly (decrease by vs increase by ) decrease the pacing to reclaim the available bandwidth. A sudden decrease is avoided to account for imperfect estimates. After pacing is applied, we empirically observe that we maintain a stable throughput (orders received per second) at the ME, despite being confronted by bursts of orders. We also observe that pacing significantly reduces packet drops. In §8.7, we show pacing can achieve a reasonable balance between throughput and packet losses. Our attempts to further improve this simple pacing heuristic were not fruitful because of the noise inherent in calculating as a cloud tenant. It is possible that there are better pacing schemes that Pareto dominate our scheme on throughput and loss; for now, though, we have found our simple scheme to suffice.
Parameter Selection We pick and the increase and decrease amounts offline by running a control experiment where VMs communicate at a fixed rate which is expected to be the equilibrium rate for order submission.
5.2 Limit Order Queue (LOQ)
LOQ implements a distributed scheduling policy across all proxies of the overlay tree, improving ME’s order matching rate and reducing order latency. Its benefits are pronounced when packet queues form at the proxies.
LOQ leverages the observation that orders with prices far from the mid-price tend to wait in the LOB, even if they arrive earlier than orders closer to the mid-price. These noncritical orders can afford to remain in proxy queues without impacting the matching process. If a proxy knows the mid-price, it can prioritize outgoing orders based on their proximity to this price. Critical orders—those closer to the mid-price and more likely to be matched—are relayed to the ME first. This prioritization boosts ME throughput (orders matched per second) and reduces the latency of matched orders.
The above approach to LOQ relies on proxies having access to mid-price information. One way to provide this information is for the ME to periodically multicast the mid-price to all proxies, which would increase the ME’s workload, especially during times of high activity and order overloads. Another solution is for each proxy to maintain an approximate LOB based on the orders it relays to the ME. However, this adds the overhead of maintaining an LOB at the proxies and may diminish the benefits of LOQ if a proxy’s LOB significantly diverges from the ME’s LOB.
Mid-price agnostic LOQ: To avoid maintaining mid-price information at each proxy, we design LOQ to be agnostic of the mid-price information. A proxy needs to prioritize some orders over others based on their probability of getting matched as they arrive at ME. The distance to the mid-price is a good indicator of this probability. However, a proxy does not need the exact probability but just the ordering relationship i.e., whether the probability of one order is higher or lower than some other order that it has in the queue. We leverage this to design a mid-price agnostic LOQ as follows.
LOQ at each proxy consists of two separate priority queues, one for bids and one for asks. The bids queue sorts the orders by price in descending order while the asks queue sorts the orders by price in ascending order. This reflects the ME behavior that the highest bid and lowest ask are selected for a potential match. A proxy consults the two priority queues in a round-robin fashion for dequeuing the orders and forwarding them to ME. This design eliminates the need for mid-price information at the proxies, simplifying the system.
LOQ also needs to respect price-time-priority that an ME uses for order matching. Each priority queue at the proxies first sorts the orders by price, as discussed above, and then by order generation timestamp, . This ensures that when a proxy dequeues an order from LOQ for relaying to ME, another order with the same price and lower does not exist in the queue. A sequencer delay is used by ME to account for any network reordering: An ME only dequeues an order for processing if it was received at least ago which gives enough time, with high probability, to the other orders to arrive and get prioritized if they deserve to do so e.g., they have a lower or an LOQ deemed it to arrive at ME earlier.
5.3 Deconstructing LOQ
LOQ improves order matching rate and reduces latency of matched orders by reordering queues at the proxies. A natural question is whether we are sacrificing correctness by such reordering. The reordering only aims to prioritize orders with a higher likelihood of getting matched, i.e., those with a lower chance of remaining in the LOB and idling the ME upon arrival at the ME. Intuitively, it seems LOQ does not affect the matching behavior of ME, but we have not formally shown that. Empirically, we use inbound fairness333Inbound fairness is distinct from the (outbound) multicast fairness to quantify how much LOQ changes the behavior of ME.
Inbound fairness: For an ME, it’s critical to ensure that orders generated earlier are matched before those generated later if both provide the same market value. This concept of inbound fairness is a key feature of matching algorithms like price-time priority. Any infrastructure for exchanges needs to ensure an ME can do the above: if an order submitted earlier () should be matched before a later order () then the ME should have received both by the time it chooses for execution.
[…]

### v11

@@ 4.4. Remarks on Multicast Packet Losses
In on-premises exchanges, the multicast service does not provide reliable delivery and Jasper adopts the same (and uses UDP in the outbound communication). Instead, a separate ”rewinder” service is deployed that handles any retransmission requests from MPs [ASX, , Jane Street and Threads, ]. MPs use sequence numbers of the messages to detect gaps and request a retransmission. The exchange infrastructure is engineered to provide negligible packet losses so the MPs only have to contact the rewinder rarely. Jasper multicast service needs to similarly provide low losses. Our hedging techniques help lower the packet losses, but we observe that even without our hedging, the losses in outbound direction are small: we observe 0.007% lost packets in a benchmark on GCP while exhausting 80% of egress bandwidth (details in Appendix F).
5. Orders Submission Service
On the inbound side, we aim to provide the order submission service that can achieve both inbound fairness and efficient order matching. To do so, we introduce a sequencer at the exchange server to enable fair processing of orders. Meanwhile, we install a novel scheduling policy, namely Limit Order Queue (LOQ) policy, on each gateway, which can significantly improve the ordering matching rate of the exchange. For simplicity, we will first describe the sequencer and LOQ in a setting where gateways are directly connected to the exchange server. Then, we continue to explain how the multicast tree is integrated to substantially improve the performance of the order submission service.
5.1. Sequencer
The exchange system has to provide a fair chance of trading to all the MPs, which requires the exchange to process MPs’ orders following their generation order. To satisfy this inbound fairness requirement, on-prem exchanges equalize the latency between MP servers and the exchange server by connecting them with the cables of the same length. However, such an approach is not feasible in the public cloud where the tenants have no access to the underlying infrastructure. Therefore, we seek an alternative approach to achieving inbound fairness by using synchronized clocks, which have become available in today’s public cloud.
In Jasper, we synchronize the clocks on both MPs and the exchange VMs with an accuracy of 10s of nanoseconds using Huygens [Geng et al., 2018],555Given that we deal with latencies on the order of microseconds, such a clock synchronization accuracy is exceptionally good. a network effect based clock synchronization algorithm robust to latency variance. We place a sequencer at the exchange server, which aims to hold the incoming orders and releases the orders according to the global FIFO order i.e., the sequencer ensures that an exchange sees the orders with timestamps in the non-decreasing order. The sequencer only releases an order of an MP to the ME when (i) it sees higher timestamped orders from every other MP and the (ii) the current order has the lowest timestamp among the orders present at the sequencer. 666 The ties on the timestamps are broken based on the MP ID. For abetting liveness (i.e., sequencer does not block processing of some orders for long periods), gateways periodically generate dummy orders on behalf of inactive traders – similar to the heartbeats employed by DBO [Gupta et al., 2023]. Given synchronized clocks and ordered delivery per participant (e.g., via TCP), the sequencer ensures safety, i.e., inbound fairness. This is different from the prior works (§7): (i) CloudEx violates safety even with accurate clock synchronization because it waits for orders only until a set timeout and, (ii) DBO holds safety for only a subset of trades (i.e., trades that only depend on specific market data points), although it does not assume synchronized clocks. Jasper’s safety comes at the cost of liveness as a failure of a gateway will block the sequencer. As the common practice in distributed systems [Lamport et al., 2009, Choi et al., 2023, cma, , Apache Software Foundation, ], the failure detection of gateways is be conducted by a standalone configuration manager (e.g., Zookeeper), which periodically checks the health of each gateway and reconfigures the membership for the Jasper cluster. We omit the details of the configuration manager since it is beyond the scope of this work. The sequencer protocol runs as follows.
5.1.1. Sequencer Protocol
There are MP VMs where denotes MP. Each is connected to the exchange server via a reliable channel (e.g., TCP channel) that provides in-order delivery. Clocks of MPs and the exchange server, where the sequencer runs, are synchronized. MPs generate order messages that are sent to the exchange server. The exchange server maintains a limit order book (LOB) and runs continuous ¡price, time¿ priority matching algorithm on all the incoming trade orders. We use to represent an order message from with an order generation timestamp . At the exchange server, messages are fed to the sequencer. A sequencer works in a streaming fashion and takes the messages as input and releases them to the ME. Before releasing a message, it is sequenced as described in the following.
The sequencer maintains where is a priority queue with lexicographic ordering on ¡, ¿ of messages. The sequencer supports enqueue and dequeue operations.
Enqueue: As a new message from a arrives at the sequencer, (i) it is added to and, (ii) if now contains non-zero number of messages from each , dequeue is invoked.
Dequeue: One message is dequeued from . is considered sequenced at this point and presented to the ME.
The above sequencer enqueue and dequeue operations provide safety at the cost of degraded liveness. For abetting liveness, each gateway generates a dummy message after every period time units on behalf the corresponding if has not generated a message during the last period time units. The sequencer dequeue discards a instead of presenting to the ME if is a dummy message where each message contains a field representing whether it is a dummy.
5.1.2. Implementation
Algorithm 1 presents an efficient implementation of the sequencer’s enqueue and dequeue routines. Enqueue is made lightweight so that incoming messages can be processed as quickly as possible and a queue formation at the ingress of the exchange’s VM can be avoided. Enqueue is invoked at each new message while Dequeue runs in a separate thread indefinitely.
5.2. Limit Order Queue
As the common practice in today’s exchange systems, we employ TCP to provide reliable delivery of orders (from MP to exchange). During periods of bursty activity in the market, incast congestion at the exchange server’s ingress occurs, leading to retransmissions which in turn leads to queue build-ups at the gateways (because of backpressure). To tackle the queuing delays incurred at the gateways, we develop a scheduling scheme –Limit Order Queue (LOQ) scheduling–to schedule the orders at each gateway, which can effectively reduce order matching latency and improve the order matching rate at ME during periods of bursts.
5.2.1. LOQ protocol
At a high level, our scheduling scheme installs a special queue (LOQ) at each gateway. LOQ takes the orders as input and schedules them in a way that keeps the matching engine from idling; ME idles if it receives orders that cannot be matched and just needs to be put into the limit order book. Specifically, LOQ categorizes the orders into two classes: In an exchange, orders closer to the mid price – critical orders – gets matched by the matching engine before the other orders – non-critical orders – which reside in the LOB and wait for the mid price to move closer. LOQ leverages this domain knowledge to identify and prioritize critical orders over non-critical orders so that the matching engine does not waste time in processing orders that are not going to get matched and only needs to be put in the limit order book while the critical orders remain to be processed.
LOQ scheduling requires two parameters: (i) mid-price and, (ii) action window . As gateways receive all the market data, they have enough information to infer . Also note that by design of our outbound communication, all the gateways receive each market datum at the same time (with high probability) so all the gateways simultaneously have enough information to infer . Here the action window indicates irrationality tolerance: (i) if an asset is available for purchase for a price then a buyer can bid on it with a maximum price of and, (ii) if an asset has a highest bid of price then a seller can ask for a minimum price of . Without loss of generality, we can replace by . The parameter is configured by the exchange operator.
An incoming order is categorized as critical if it has a price in the range [, ], otherwise it is non-critical i.e., if a bid has a price in the range (, ) or an ask has a price in the range (, ), then it is non-critical.
An LOQ is a priority queue that sorts the orders lexicographically by tuple ¡, , ¿ where starts from 0 and is incremented every time changes, is 0 if an order is critical otherwise it is 1 and denotes the generation timestamp of an order. LOQ maintains the global variable while is calculated per order while enqueuing it. LOQ runs in the gateways where every order generated by the trading algorithm is enqueued to it and a separate thread running the dequeues forwards the orders to the exchange server. The design of the priority function is motivated by the need to not impact inbound fairness as we explain in the following.
5.2.2. Inbound fairness under LOQ
For safety, sequencer assumes that messages of a single MP arrive at the sequencer in their generation order. As LOQ at a gateway reorders the messages, this assumption is violated. However, we claim that fairness is still achieved as we have designed the LOQ policy to do so. Here we provide the intuition of our claim while Appendix I contains a detailed intuitive proof.
Assume there is a static mid-price. Instead of looking at the order in which messages/orders are presented to the ME, let’s look at the order in which the orders are executed by the ME. By design LOQ ensures that if an order is selected for execution by ME, then all the older executable orders must have been executed already. Executable orders are the ones whose price falls in [, ] i.e., the critical orders and LOQ ensures that orders falling in this range are sorted by their timestamps (and a sequencer consults such LOQ instances and sequences them in order of timestamps). So, ME never executes a critical order if it has not executed all the older critical orders which constitutes fairness. When the mid-price moves, LOQ ensures that the messages generated after the movement have lower priority than all the prior messages (due to being the first element of the tuple used by LOQ’s priority function), leading to fairness. We conclude this section by reiterating the assumptions that are needed: (i) clock synchronization and (ii) all MPs simultaneously inferring the mid-price movement. Both assumptions are valid in practice with high probability.
5.3. Reusing Multicast Tree For Scaling Order Submission Service
As the number of MPs grows, the exchange’s performance degrades as it cannot keep up with offered load. However, we find that even if the cumulative offered load to an exchange stays constant, the growing number of MPs leads to performance degradation because of the increase in instantaneous load: The simultaneous order submission from a large number of MPs overwhelms the exchange’s ingress. For scaling, the exchange systems usually adopt a sharding approach by partitioning the assets across multiple exchange servers [Ghalayini et al., 2021]. We focus on enhancing the performance of a single exchange server/shard. Our technique is orthogonal to the partitioning based scaling of the exchange.
With a large number of MPs simultaneously submitting orders to the exchange server, we observe a significant packet loss that leads to decrease in the exchange’s throughput and increased latency for orders. We employ a simple strategy: reuse the multicast tree in the reverse direction for proxying the MPs messages to the exchange. The reduced fan-in at the exchange server reduces the packet drops while packet queues form at the tree nodes. We run LOQ at each tree node to service the queues, leading to a significantly high order matching rate of the ME and low latency for orders.
Furthermore, running LOQ at each proxy node leads to better scheduling compared to running LOQ at just the gateways. A proxy node has messages from several MPs in its queue and can prioritize messages of one MP over the messages of another MP –as long as it does not affect fairness– where an LOQ at a gateway only has the opportunity of prioritizing messages of an MP over the messages of the same MP. TCP connections are terminated at each tree node to allow the LOQ to reorder the messages.
Fairness when using a tree. Achieving fairness requires that an LOQ instance at each tree node is accompanied by a sequencer instance so that LOQ can operate on the messages of all children nodes fairly. Without a sequencer, a delayed message of an MP may not get assigned its proper priority by LOQ, which would eventually result into safety (fairness) violation at the exchange server. When a sequencer is used at every tree node, it only accounts for messages from its children and not all MPs, so the parameter changes accordingly in Algorithm 1. At each tree node, sequencer discards the incoming dummy messages and each proxy node generates new dummy messages.
Remarks on other methods for dealing with incast: Several methods [Montazeri et al., 2018, Cho et al., 2023, 2020, Alizadeh et al., 2010] for incast mitigation and overload control have been proposed in the literature. The methods are largely orthogonal to our technique of utilizing the tree in reverse which we can do solely because we have the control of an application that lends us a tree of VMs. Credit-based overload control schemes [Cho et al., 2023, 2020] are composable to Jasper, enhancing Jasper’s performance by reducing packet losses/retransmissions. However, a tree further gives us the opportunity to do better scheduling (LOQ) on multiple clients (MPs) data which is not possible without a tree/intermediate nodes and forming the queues only at the order gateways; as would be the case if we use existing techniques.
6. Evaluation
We focus on comparison against a prior system, CloudEx [Ghalayini et al., 2021]. As Jasper can be viewed as CloudEx augmented with techniques to enhance performance under large number of participants, the comparison with CloudEx shows how much CloudEx can be enhanced. We also present an ablation study as Jasper comprises of several techniques put together. System name Jasper and Onyx is used interchangeably in this tech report. We answer the following questions.
How does Jasper multicast latency compare with AWS TGW-based multicast? §6.1
How much outbound fairness is achieved? §6.2
How does multicast fairness and performance compare to CloudEx? §6.2
How scalable is Jasper? §6.3
How does order submission compare to CloudEx? §6.4
How much does LOQ help in enhancing ME’s throughput and lowering orders’ latency? §9
- (7)
How can Jasper help DBO? §6.6
For most of the experiments, we use 100 MPs and a 5K multicast messages per second (MPS) rate. Each MP has a separate VM. For showing scale, some experimemts utilize 1000 MPs and is mentioned in the respective sections. No two MPs share a VM. In practice, multiple MPs could be hosted in a single VM to further enhance scalability. During our evaluation, we define two metrics to quantify the performance of Jasper and the baseline systems. The overall multicast latency (OML) refers to the latency experienced by the last receiver that receives a multicast message. The delivery window size (DWS) is the maximum difference in the latency of any two receivers. We use the Huygens algorithm [Geng et al., 2018] that synchronizes clocks of all the VMs with a offset of .
6.1. Multicast Latency Comparison
AWS Transit Gateway: AWS-TGW-based multicast is provided by AWS [Services, ]. It requires the sender to send message to a gateway which then replicates and sends one copy to each receiver. It can support at most 100 receivers.
Figure 6 shows that Jasper distinctly outperforms DU and AWS-TG. This experiment was performed entirely on AWS. The median latency for Jasper is which is lower than the latency of AWS TG () and lower than the latency by DU (). At 90p, Jasper shows lower latency than both DU and AWS TG. Moreover, Jasper shows predictable latency as higher percentiles are very close to the median in contrast to the other techniques.
6.2. Outbound Fairness Comparison
We measure the DWS of Jasper and CloudEx to compare their outbound fairness performance. A smaller DWS indicates a higher level of outbound fairness (i.e., simultaneous delivery).
Figure 6(a) shows that Jasper can achieve a DWS of at very high percentiles (up to ). Without proxy hedging, the DWS becomes larger at earlier percentile (). CloudEx achieves fairness but at the cost of high OML.
Figure 6(b) shows that OML increases significantly for CloudEx and Jasper with no hedging. In these systems, the deadlines calculated by the hold-and-release are far into the future to cover the high latency variance. This leads to high holding duration at the receivers for CloudEx and Jasper without hedging as shown in figure 6(c), increasing OML.
We do not include a comparison with DBO here as DBO changes the definitions of fairness to not rely on simultaneous delivery. However, the latency of DBO’s market data multicast would be slightly worse than a Direct Unicast (DU) approach discussed in §6.1 as it does batching and pacing of multicast packets. Jasper outperforms DU.
6.3. Scaling Jasper Multicast
Scaling N to 1000. In contrast to the previous fair multicast solutions that only work with a few 10s of receivers (e.g., [Gupta et al., 2023, Gong et al., 2024, Services, ]), Jasper aims to implement a more scalable multicast service which can support 1000 receivers and potentially even more. Figure 8 plots Jasper’s multicast latency for 100 receivers and 1000 receivers. Our overlay techniques for reducing latency enable Jasper to keep a graceful latency growth as the number of receivers increase.
Fairness for an increasing number of receivers. When employing hold-and-release, a median DWS of is achieved for . However, the highest percentile at which DWS stays decreases as increases. We define the probability of fairness (()) as the highest percentile where DWS is . Table 1 shows OML and () for = 100 and 1000.
Throughput. Jasper can utilize 80% of the egress bandwidth of a c2d-highcpu-8 VM on GCP while maintaining negligible packet losses. The specified VM has an egress bandwidth of 16Gbps. With 466B packets, we achieve a multicast message rate of 350K MPS, without proxy hedging. The OML stays the same as we increase the message rate from 5K to 350K, but increases . In the experiment, we have 100 receivers and 10 proxies, with no proxy hedging. The sender and each proxy replicate each message times hence, with 350K messages per second, the sender generates 3.5M packets per second, utilizing 81.5% of the available egress. The proxy hedging would reduce the multicast rate according to the factor as each proxy replicates each multicast message times, e.g., with , we achieve 175K messages per second.
6.4. Orders Submission Performance
Figure 9 shows the order matching rate of the exchange server for three systems. Onyx outperforms CloudEx by atleast an order of magnitude. 100 MPs cumulatively generate 100K orders per second each where periodic bursts occur (shown as shaded regions). The order generation rate becomes during the bursts. MPs stop generating the orders after 20s. Onyx is able to keep up with the offered load and finishes processing the orders right after the MPs stop. Onyx achieves higher order matching rate when bursts occurs due to LOQ. CloudEx achieves significantly low order matching rate and builds up queues that are processed even after the MPs stop. The median latency of orders with Onyx is lower compared to CloudEx.
Limit Order Queue (LOQ) Performance We compare LOQ to a first-in-first-out queue, which we refer to as SimplePQ. Unlike LOQ, SimplePQ does not differentiate between critical and non-critical orders. In this experiment, we have 10 market participants (MPs), each submitting 7K orders per second. Two intermediate proxies, running LOQ or SimplePQ, relay orders from the MPs to the ME. Order bursts occur, with each MP doubling their order submission rate every . Bursts lead to queue build ups at the proxies, where LOQ/SimplePQ operate. As shown in Figure 11, LOQ outperforms SimplePQ in ME’s order matching rate and matched orders’ latency. Thin lines represent the orders received per second by ME which are similar (but noisy) for both SimplePQ and LOQ. Thick lines represent order matching rate and orders latency in their respective plots showing LOQ’s better performance.
The advantages of LOQ stem from the presence of non-critical orders, which are kept in the gateways and proxies’ queues longer, allowing the available egress to be used for critical orders. We examine the benefits of LOQ under different ratios of critical to non-critical orders. In this experiment, MPs uniformly sample bid/ask prices from a predefined range to generate orders. We vary the portion of the bid price range that overlaps with the ask price range. These overlapping segments result in orders that get matched at ME, while other orders remain in the LOB. Figure 11 shows that, as the ratio of overlapping price levels to the total number of levels decreases, the benefit of LOQ increases. We plot the metrics for the bursts duration which are calculated from the orders matched during the bursts that we inject. A ratio of indicates that there are total price levels from which MPs sample prices to generate orders, but only price levels lead to immediate matches at ME. As increases while remains constant, the number of critical orders decreases. LOQ identifies critical orders and prioritizes them for processing by the ME over non-criticals. Consequently, LOQ demonstrates a 150% increase in ME throughput and a 99.9% reduction in latency for matched orders when .777 is commonly used in the literature [Aitken et al., 2007] We repeat one experiment and increase MPs from 10 to 1000 with and observe that LOQ distinctly outperforms SimplePQ: 98% decrease in the latency and 10% (86%) increase in overall (during bursts) matching rate.
Using Proxy Tree for Orders Submission: To conduct an ablation study of the proxy tree’s benefit in the order submission, we run an experiment without LOQ and sequencer, and measure the number of packets received per second by the exchange server with and without a tree. Figure 12 plots the throughput achieved with 100 MPs. 100 MPs send at the cumulative base rate of 100K packets/second, but create periodic bursts that reach higher than the base rate. By using the proxy tree, Jasper can improves the overall throughput by on average and by during the bursty periods.
Sequencing Overhead: We perform an experiment with 100 MPs (without LOQ) to study the throughput with and without our sequencer. We use the same load generator as above and notice an overall 25% decrease in the throughput of the exchange when using the sequencer. The throughput decrease is traded for improving inbound fairness. On the other hand, our LOQ protocol improves the Jasper throughput by assigning priorities to the packets and facilitate the order matching at the exchange. As a result, the incorporation of the proxy tree and LOQ protocol compensates for the sequencer overheads and enables Jasper to outperform CloudEx (Figure 9).
Remarks on inbound fairness: By design, our sequencer (w/o LOQ) ensures 100% inbound fairness. One of the assumptions for LOQ to respect fairness is that all MPs/gateways infer the mid-price movement at the same time which is provided by outbound fairness with a high probability. Therefore, the probability of achieving inbound fairness while employing LOQ can only be as high as that of outbound fairness which, we show in §6.2 and §6.3, is achieved by 90% probability. In §6.5.3, we will continue to show the probability of outbound fairness can be optimized to 99.9%, which is much better than CloudEx.
| N | P(F) (%) | OML |
| 100 | 92 | 175 |
| 1000 | 89 | 238 |
6.5. Outbound Communication Techniques
6.5.1. Proxy Hedging
Reduced OWDs for each receiver. Proxy hedging improves OML as the OWD to each receiver is reduced. Figure 8 shows OWD for each receiver with () and without hedging.
6.5.2. Round Robin Packet Spraying
We evaluate the impact of RRPS on the multicast latency. We utilize a base multicast message rate of 10K MPS, 125 receivers and a burst of occur every . During bursts, the message rate becomes . Table 3 shows ulticast latency decreases when RRPS is used. On GCP, we see an OML reduction of , however, on AWS the reduction can approach . With low message rates, the reduction is not significant: with message rates MPS. This happens because the inter-packet duration is small enough that any transient latency spike is not able to impact consecutive messages so RRPS does not help much by re-routing messages.
RRPS show improvements in OML when message bursts are introduced as it can distribute the bursts among several network paths. Without bursts, the OML reduction still happens but is less than on both AWS and GCP.
6.5.3. Receiver Hedging
| Setup | On AWS | On GCP | ||
| 50p | 90p | 50p | 90p | |
| Improve. (%) | 15.12 | 69.35 | 9.83 | 9.18 |
| RH | OML () | ||
| 50 | 99 | ||
| 100 | No | 139 | 248 |
| 200 | No | 141 | 243 |
| 100 | Yes | 99 | 146 |
6.6. Onyx and DBO
DBO is a recent cloud exchange that achieves fairness regardless of latency fluctuations and without clock synchronization. It does so by proposing a new fairness metric that prioritizes orders based on the response time of traders (i.e., the time to make a trade in response to a piece of market data), rather than the time at which orders were submitted. Onyx can be viewed as a network layer that is complementary to DBO’s new semantics at the application layer. In particular, Onyx can help DBO scale out to a large number of participants, particularly in the following three ways. One, Onyx’s low-latency and scalable multicast tree (and hedging) can help with disseminating market data in DBO. Two, Onyx’s use of a tree in reverse for inbound order submission can also help DBO deal with incast-type conditions during order submissions from a large number of participants. Three, because DBO does not enforce simultaneous delivery of data, DBO has an extra constraint that no MPs should directly or indirectly talk to each other as it can violate DBO’s fairness guarantees. By leveraging Onyx’s simultaneous multicast in DBO, this constraint can be lifted.
7. Related Work
We have already discussed the recent exchanges (CloudEx [Ghalayini et al., 2021], DBO [Gupta et al., 2023]) and overload control schemes [Cho et al., 2023, 2020, Montazeri et al., 2018].
Multicast: Prior works on application level multicast [Banerjee et al., 2002, Rouskas and Baldine, 1997, Parsa et al., 1998, hua Chu et al., 2002, Mokhtarian and Jacobsen, 2015] mainly focus on finding optimal paths in a network (or overlay mesh) using cost models for network links that capture heterogeneous link bandwidth or latency characteristics. Jasper focuses on (i) minimizing the delays incurred at hosts while transmitting multiple message copies, and (ii) achieving a small latency difference across receivers. On the other hand, switch-based multicast [Prabhakar et al., 1997, Guo and Chang, 1998] is not available to cloud tenants due to scalability issues [Shahbaz et al., 2019].
Collective Communication: Collective communication [Zhuang et al., 2021, Zhao and Krishnamurthy, 2023, Zhao et al., 2023] uses overlay trees for broadcast and all-reduce [Patarasuk and Yuan, 2009], but typically supports fewer than 100 receivers. For example, Hoplite [Zhuang et al., 2021] optimizes bandwidth for 10s of nodes using a tree. Jasper focuses on minimizing latency and variance, handling bursts, and achieving fairness at scale while not having an aggregation mechanism for trade orders.
8. Conclusion
This paper introduces Jasper, which provides networking support for scalable financial exchanges. Jasper systematically optimizes both inbound (order submission) and outbound (market data delivery) workflows of the exchange system. Our evaluation shows that Jasper outperforms a prior system CloudEx, in terms of fairness, throughput, and latency as well as outperforms AWS TGW-based multicast. This work does not raise any ethical issues.
Appendix A Deciding D and F for multicast tree
To understand how the end-to-end latency varies as and grow, we conduct a series of experiments with various numbers of receivers (). Table 4 displays the latencies for different configurations of ¡¿ for a given number () of receivers. We see that the latency reaches a minimum point at different depths for different values of . At a small scale (), increasing does not bring latency benefits. As the scale grows from to , the benefits of increasing while reducing grow because the reduced message replication delay and transmission delay outweigh the added overhead of new hops in the path of messages. However, as goes beyond a certain threshold (e.g., grows larger than 3 when ), the latency no longer decreases but gets worse. Based on our experiments, we find fixing usually leads to a desirable to generate a multicast tree with low latency. We establish our heuristic rule to construct the multicast tree as follows: Given receivers, we fix and then derive (round to the nearest integer). We also tested a more sophisticated alternative mechanism described in [Venkataraman et al., 2016], and observe no significant performance boost compared to our heuristic.
In our testing, we find that the variance in cloud-host VM performance [Sachidananda, ] imposes the great challenge to differentiate the “optimal” values of and from values picked using the aforementioned heuristic. We observe through repeated experimentation that minor changes in and do not show a significant performance difference with high statistical confidence. So it is sufficient to select and in the neighborhood of the optimal value, which is why our heuristic is effective. We further find that a unit increment in comes at the added latency of and a unit increment in adds per layer.888Using our high-performance implementation on a c5.2xlarge VM in AWS. Message size is 460B. A tree constructed using linear models based on these unit increments performs comparably to the tree constructed using our heuristic of maintaining close to 10 and . An example scenario for deciding tree structure: For , , would be 2. For , , would still be 2 and F would need to adjusted accordingly to support all 200 receivers. So F would come around to be be 14 as .
Appendix B Scalable Simultaneous Delivery
Financial trading needs to ensure the fairness of the competition. Fairness in data delivery [Ghalayini et al., 2021, Gong et al., 2024] means that the market data from the exchange server should be delivered to every MP at the same time so that an MP may not gain an advantage over the others during the competition. While there are also some recent works trying to alter the definition of fairness [Gupta et al., 2023, Mavroudis and Melton, 2019], these variant definitions require more research before they are confidently adopted by the exchanges. Therefore, Jasper uses the original definition of fairness employed by on-premises financial exchanges. In short, in Jasper the fair delivery of data means simultaneous delivery of market data to all the multicast receivers.
Realizing perfect simultaneous data delivery to multiple receivers is theoretically unattainable [Gmytrasiewicz and Durfee, 1992]. Nevertheless, Jasper tries to empirically minimize the spatial (i.e., across receivers) variance of the latency of messages. Our hedging design (§4.2) has created favorable conditions to minimize the spatial variance. By using hedging, Jasper can achieve consistently low variance for each receiver over time, and the spatial latency variance across receivers is kept low. Beyond this, we employ a hold-and-release mechanism (by using synchronized clocks) to eliminate the residual spatial variance at the end hosts and enforce simultaneous delivery across receivers. The hold-and-release mechanism was introduced by CloudEx [Ghalayini et al., 2021] but it does not scale well and leads to high end-to-end latency (§6.2). We describe a modified hold-and-release mechanism that helps us scale further, while maintaining a low end-to-end multicast latency.
Hold & Release mechanism. To implement the hold-and-release mechanism, Jasper leverages the accurate clock synchronization algorithm, Huygens [Geng et al., 2018], to synchronize the clocks among the sender and receivers. Receivers keep track of the one-way delay (OWD) of the messages received from the sender. Each receiver takes the 95th percentile of its OWD records at regular intervals and sends the results back to the sender using an all-reduce mechanism explained later. The sender calculates the maximum of these OWDs called Global OWD for messages using the gathered statistics: where is the OWD estimate reported by the -th multicast receiver.
Once an has been calculated by the multicast sender, it attaches deadlines to all outgoing messages. A deadline is calculated by adding to the current timestamp when sending a message. Upon receiving a message, a receiver does not process it until the current time is equal to (or exceeds) the message’s deadline. This mechanism leads to almost simultaneous delivery, modulo clock sync error. In §G, we discussed possible security mechanisms to ensure that a receiver (an MP) waits until its deadline to process a message, when it’s in the MP’s self interest not to wait.
[…]